Excerpt taken and modified from Government of Pakistan Investment bond’s Auction result:

Auction for 3, 5, 10- and 20-year Pakistan Investment Bonds was held on March 20, 2019 with settlement date on March 21, 2019. The coupon rates for 3, 5, 10- and 20-year bonds are 7.25%, 8.00%, 8.75%, and 10.75% respectively. The accepted bid summary and result is as under:


a. Using weighted average yield % p.a calculate the following rates:
a. The 2-year rate 3 years from now
b. The 5-year rate 5 years from now
Given the following information what do you think will be the shape of the Pakistan investment bond yield curve. Explain support your answer with the help of the latest SBP Monetary policy report.
1
Expert's answer
2020-12-01T05:04:20-0500


a.b,"\frac{1\times7.25+2*\times7.25+3\times7.25}{3}=14.5"

"\frac{1\times8.0+2\times8.0+3\times8.0+4\times8.0+5\times8.0}{5}=24"


b. At its meeting on 23rd November 2020, the Monetary Policy Committee (MPC) decided to maintain the policy rate at 7 percent. However, these supply-side pressures are likely to be temporary and average inflation is expected to fall within the previously announced range of 7-9 percent. Bond yields depend directly on the Central Bank rate. This means that the bond yield directly depends on the Central Bank's rate will be in the range of 7-9 in 21 years%